OpenStudy (anonymous):

perimeter of parallelogram is 48 in. AB is 5 in less than BC ( use the given information to find the lengths of all four sides of parallelogram ABCD)

OpenStudy (radar):

Let x = shorter side of parallelgram Let x+5 = longer side of parallgram 2(x) + 2 (x+5) = 48 4x = 48 -10 = 38 x = 9.5 x+5 =14.5 sides are 9 1/2 and 14 1/2

OpenStudy (anonymous):

Lovely help, thankyou radar!